OpenStudy (anonymous):

In 1994, the average price of a new domestic car was $16,930. In 2002, the average price was $19,126. Based on a linear model, what would be the approximate predicted average price for 2008? Question 8 answers $23,000 $21,300 $20,800 $18,600

OpenStudy (anonymous):

DO the slope : y2-y1 19126-16930 X-19126 m=------ = -------------- =274.5= --------------- x2-x1 2002-1994 2008-2002 X=274.5(2008-2002)+19126 = 20,773 or 20,800 answer

OpenStudy (anonymous):

thanks boo (:

OpenStudy (anonymous):

wc

OpenStudy (anonymous):

Y2-Y1 hope you understand the slope m = -------- X2-X1

OpenStudy (anonymous):

hi bailey the equation on your next question was missing the X..i mean the prob before this one..about investment